# Natural-artificial hybrid swarm: Cyborg-insect group navigation in unknown obstructed soft terrain

### v2

@@ Discussion
Third, neighbors can aid a cyborg in recovering from an overturn. In the presence of neighbors, the overturned cyborg can use neighbors as supporting points for more efficient uprighting. Fig. 7a illustrates three cases of self-attempted and neighbor-aided recovery from overturns. The discrepancy in recovery performance is particularly evident in Case 3, where a cyborg struggled for 26.5s before grasping a passing neighbor and then successfully recovered in only 4.5s. Snapshots of Cases 1 and 3, where cyborgs recovered with the help of neighbors, are provided in Figs. 7b and 7c, respectively. The proposed TGI control algorithm ensures the presence of neighbors around each cyborg and, in turn, enhances the robustness of the multi-cyborg system.
Figure 7: Recovery from an overturn. a Time for a cyborg to recover from an overturn for three trials of experiments. The grey bar denotes the time a cyborg attempted to recover by itself. The bar’s bottom and top indicate the start and ending times. Similarly, the purple bar denotes the time taken with help from a neighbor. In Cases 1 and 3, the cyborg in the overturned state first attempted and failed to recover by itself for a while, then successfully recovered by grabbing a passing neighbor. b The snapshots of the recovery process of Case 1: the overturned cyborg grabbed an approaching neighbor and, with its help, successfully recovered. c The snapshots of the recovery process of Case 3. See Movie S5.
In summary, this paper studied a swarm navigation problem for a multi-cyborg system. A control strategy has been proposed to tackle this problem and verified under experiments. The proposed control algorithm offers several notable advantages. Firstly, it efficiently utilizes the instincts of insects, exploiting their inherent adaptability to various environments. Additionally, by leveraging these instincts, the algorithm reduces the frequency of electric stimulations, consequently extending the operational duration of the system. Moreover, the decentralized nature of the control algorithm ensures its scalability to a larger swarm, as each cyborg makes decisions based on only local information.
Note that although the control algorithm proposed in this work is theoretically fully decentralized, the cyborgs in our experiments were provided data from a centralized motion capture system. We are considering more advanced positioning systems to address this issue in future work. A feasible solution is using a micro-nano inertial measurement unit (i.e., the Honeywell MV 60 high-precision micrometer) and radio frequency identification (RFID) technology for localization. The advantages of this solution are as follows: firstly, the micro-nano inertial navigation unit can match the body size of cockroaches and provide centimeter-level positioning accuracy within a certain period; secondly, to address the error accumulation property of accelerometers, RFID modules can be deployed around the site which update the location information of cyborgs at intervals, and thus zero out the accumulated errors of the accelerometers. This solution can be a promising direction for the practical utilization of the proposed control strategy.
@@ Acknowledgements
Funding: The authors received funding from JST (Moonshot R&D Program, Grant Number JPMJMS223A). Author contributions: N. W., H. S., and M. O. conceived and designed the research. Y. B., M. O., and N. W. developed the swarm control algorithm. P. T. T. N., H. D. N., D. L. L., and K. K. built the experimental platform. Y. B., P. T. T. N., Q. H. H., K. K., and Y. X. S. T. prepared the cyborg insects. Y. B., P. T. T. N., H. D. N., Q. H. H, Y. D., and Y. X. S. T. collected the raw data for the cyborg experiments. Y. B., J. S., and Y. D. conducted the data analysis. Y. B., J. S., P. T. T. N., M. O., N. W., and H. S. contributed to the preparation of the manuscript. N. W., H. S., and M. O. supervised the research. Competing interests: The authors declare no competing interests. Data and materials availability: The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Supplementary materials
Materials and methods
Fig. S1
Movies S1 to S5
References (33, 34)

### v3

@@ Discussion
Third, neighbors can aid a overturned cyborg for more efficient uprighting. Although neighboring cyborgs may navigate around an overturned cyborg in a manner similar to how they avoid obstacles, they also play a passive yet crucial role in facilitating its recovery. The overturned cyborg can utilize interactions with nearby teammates, such as physical contact or grasping, to aid in up-righting itself. This cooperative interaction enhances the overall resilience of the swarm without requiring direct, algorithmic intervention. The bar chart in Fig. 5b illustrates three cases of self-attempted and neighbor-aided recovery from overturns. The discrepancy in recovery performance is particularly evident in Case 3, where a cyborg struggled for 26.5s before grasping a passing neighbor and then successfully recovered in only 4.5s. Snapshots of Cases 1 and 3, where cyborgs recovered with the help of neighbors, are also provided in Fig. 5b, respectively. The proposed TGI control algorithm ensures the presence of neighbors around each cyborg and, in turn, enhances the robustness of the multi-cyborg system.
In summary, our approach offers advantages over conventional control methods in terms of safety, endurance, and adaptability when applied to cyborg swarms. Without relying on precise locomotion control, our algorithm achieves a compact swarm through adjustable parameter design and effectively harnesses insects’ instincts for efficient obstacle negotiation in complex, obstructed terrains. Additionally, our algorithm mitigates the risk of habituation and extends operational time by reducing the need for control inputs by an average of $50\%$ compared to the individual control. This reduction is accomplished through the free-motion and move-toward-crowd rules, which leverage the insects’ natural motion while maintaining swarm cohesion. Furthermore, our algorithm mitigates the risk of entanglements that can destabilize the swarm and damage the cyborgs, using a switching mechanism between stimulated and free motions based on local swarm compactness. Experiments demonstrate that this design reduces the number of entanglements by over $85\%$ on average compared to the BOIDS method, enhancing the overall safety of the swarm system.
Note that although the control algorithm proposed in this work is theoretically decentralized, the cyborgs in our experiments were provided data from a centralized motion capture system. To achieve decentralized sensing, enabling cyborgs to directly measure distances between each other, we propose a hybrid approach for future work that combines Inertial Measurement Unit (IMU) and Ultra-Wideband (UWB) technologies. Specifically, each cyborg insect will be equipped with a low-power IMU to measure yaw, pitch, and roll angles. This data will be used to estimate position by analyzing body vibrations during locomotion, with speed derived from step frequency and position calculated through time integration. To mitigate IMU drift, UWB modules will be deployed to periodically update each insect’s position using precise distance measurements. This decentralized hybrid IMU-UWB system can provide robust, accurate localization for the cyborgs without a motion capture system. It can be a promising direction for the practical utilization of the proposed control strategy.
Methods
Control algorithm design
Let us define the position of a cyborg by $\textbf{q}_{i}(t)\in\mathbb{R}^{2}$, velocity by $\textbf{V}_{i}(t)\in\mathbb{R}^{2}$, and orientation by a unit vector $\textbf{o}_{i}(t)\in\mathbb{R}^{2}$, at time $t$. Without loss of generality, we assign $i=0$ for the leader and $i\neq 0$ for followers. For cyborg $i$ at time $t$, we define the sets of its neighbors by
$\mathcal{N}_{i}(t)=\{j\in\{0,\ldots,N\}\backslash\{i\}\ |\ 0<\|\textbf{q}_{j}(t)-\textbf{q}_{i}(t)\|\leq r_{s}\},$ (1)
where $N$ is a positive integer and $r_{s}$ is a positive constant, denoting the radius of the sensing range. Similarly, we define
$\mathcal{M}_{i}(t)=\{j\in\{0,\ldots,N\}\backslash\{i\}\ |\ 0<\|\textbf{q}_{j}(t)-\textbf{q}_{i}(t)\|\leq r_{\textrm{free}}\},$ (2)
where positive constant $r_{\textrm{free}}$ denotes the radius of the free range. Thus
$m_{i}=|\mathcal{M}_{i}|,$ (3)
is the number of neighbors in the free range of cyborg $i$. Note that $r_{\textrm{free}}<r_{s}$.
Based on these definitions, the swarm control algorithm is detailed as follows. It consists of two rules:
  • FM rule: for cyborg $i\in\{1,\ldots,N\}$ (followers), if the number of neighbors within the free range $m_{i}>M$ where $M$ is a positive integer, it engages in free motions; otherwise, it follows the rule below.
  • MTC rule: for cyborg $i\in\{1,\ldots,N\}$, its sensing range is divided into s sectors, and steered to the target sector (the one with the largest number of neighbors where the leader is counted as $N$ cyborgs). On the other hand, the leader $i=0$ is driven to the goal position. The aforestated functions are realized by
    $\textbf{f}_{m,i}(t)=\begin{cases}\textbf{p}^{\textrm{goal}}(t)-\textbf{q}_{i}(t),&i=0,\\ \textbf{p}^{\textrm{target}}(t)-\textbf{q}_{i}(t),&i\in\{1,\ldots,N\},\end{cases}$ (4)
    where $\textbf{p}^{\textrm{goal}}(t)$ and $\textbf{p}^{\textrm{target}}(t)$ denotes the position of a target cyborg (an arbitrary cyborg in the target sector) and the goal for the leader, respectively. Then, the desired orientation of cyborg $i$ can be computed by
    $\textbf{o}_{i}^{\textrm{des}}(t)=\mathbf{\phi}(\textbf{f}_{m,i}(t)),$ (5)
    where $\phi:\mathbb{R}^{2}\backslash\{\textbf{0}\}\rightarrow\mathbb{R}^{2}\backslash\{\textbf{0}\}:\textbf{x}\mapsto\frac{\textbf{x}}{\|\textbf{x}\|}.$
    Note that the following repulsion function can also be integrated into the MTC rule to achieve a rather sparse swarm: cyborg $i$ is steered away from a neighbor $j$ at time $t$, according to
    $\textbf{f}_{r,i}(t)=\begin{cases}\textbf{0},&i=0,\\ -\sum_{j\in\mathcal{N}_{i}(t)}\frac{1}{\|\textbf{q}_{j}(t)-\textbf{q}_{i}(t)\|^{2}}(\textbf{q}_{j}(t)-\textbf{q}_{i}(t)),&i\in\{1,\ldots,N\}.\end{cases}$ (6)
    Combining these functions, the desired velocity of follower $i$ is computed according to
    $\textbf{V}_{i}^{\textrm{des}}(t)=k_{m}\textbf{f}_{m,i}(t)+k_{r}\textbf{f}_{r,i}(t),$ (7)
    where $k_{m}$, $k_{r}$ are constant weights. Then, the desired position $\textbf{q}_{i}^{\textrm{des}}(t)$ of cyborg $i$ can be computed by
    $\textbf{q}_{i}^{\textrm{des}}(t)=\textbf{q}_{i}(t)+\textbf{V}_{i}^{\textrm{des}}(t)\Delta t,$ (8)
    where $\Delta t$ denotes the time step.
Thus, one defines the position error of cyborg $i$ by $\textbf{e}_{q,i}(t)=\textbf{q}_{i}^{\textrm{des}}(t)-\textbf{q}_{i}(t)$ and the orientation error by $\textbf{e}_{o,i}(t)=\textbf{o}_{i}^{\textrm{des}}(t)-\textbf{o}_{i}(t)$. The angle between $\textbf{o}_{i}^{\textrm{des}}(t)$ and $\textbf{o}_{i}(t)$ is denoted by $\theta_{i}$, and the speed of agent $i$ is represented by $v_{i}$. The computed $\textbf{e}_{q,i}(t)$, $\textbf{e}_{o,i}(t)$ and the measured $\theta_{i}$, $v_{i}$ are passed to the trajectory tracking algorithm summarized in Algorithm 1 in the Supplementary Information.
Data availability
The data that support the findings of this study are available within the paper and Supplementary Information files. Source data are provided in this paper.
Code availability
The code that supports the findings of this study is available within the Supplementary Information files.
@@ Acknowledgements
The authors received funding from JST (Moonshot R&D Program, Grant Number JPMJMS223A).
Author contributions
@@
N. W., H. S., and M. O. conceived and designed the research. Y. B., M. O., and N. W. developed the swarm control algorithm. P. T. T. N., H. D. N., D. L. L., and K. K. built the experimental platform. Y. B., P. T. T. N., Q. H. H., K. K., and Y. X. S. T. prepared the cyborg insects. Y. B., P. T. T. N., H. D. N., Q. H. H, Y. D., and Y. X. S. T. collected the raw data for the cyborg experiments. Y. B., J. S., and Y. D. conducted the data analysis. Y. B., J. S., P. T. T. N., M. O., N. W., and H. S. contributed to the preparation of the manuscript. N. W., H. S., and M. O. supervised the research.
Competing interests
The authors declare no competing interests.
Figures
Figure 1: Overview of the cyborg swarm navigation. (a) A cyborg insect. In this work, the Madagascar hissing cockroach (Gromphadorhina portentosa) was chosen for the cyborg system. A backpack circuit board attached to the insect was designed to house the necessary systems, including the locomotion control system and wireless communication module. It was powered by a rechargeable LiPo battery. How to prepare a cyborg insect is detailed in the supplementary materials. (b) An illustration of cyborg swarm navigation (top view). A decentralized algorithm was proposed to navigate the cyborg swarm to a designated goal through a sandy area in the presence of hills and obstacles. All the cyborgs have no information about obstacles and hills in the field. (c) The side view and (d) the back view of a cyborg swarm. A comprehensive presentation of this work can be found in Supplementary Video 1.
Figure 2: Navigation algorithm design for a cyborg swarm. (a) The cyborg swarm consists of a leader and several followers. Each cyborg can detect neighbors within a limited sensing range, and distinguish the leader and followers. Only the leader is given the position of the goal. (b) An illustration of the proposed navigation algorithm. It consists of two parts: motion planning and trajectory tracking. The motion planning algorithm provides the desired positions of cyborgs for the next step based on their local information and passes it to the trajectory-tracking algorithm. The trajectory tracking algorithm computes the corresponding amplitude and types of stimulation (left, right, or acceleration) to be applied to the insects.
Figure 3: Summary of swarm navigation experiments. (a) An illustration of experiments in an obstructed soft terrain (front view). (b) Path plot of Exp 1. (c) The top view of Exp 1. The video of Exp 1 is Supplementary Video 2. (d) Path plots of Exp 2 to 10. In Exp 2, the leader’s marker was untrackable for more than 5s, but the followers maintained group cohesion by moving to the leader’s last known position. They resumed tracking the leader after it gained its trackability. (e) The degree of autonomy for the leader under conventional control (yellow) and the followers under our proposed control algorithm (blue) in 10 trials of experiments. The error bars denote the standard deviation. Our approach nearly doubles the cyborgs’ free-motion time, effectively mitigating the likelihood of habituation of insects. Source data are provided as a Source Data file.
Figure 4: Comparison between the experimental results under the conventional BOIDS and the proposed TGI algorithms. (a) The frequency of entanglements in 10 trials of experiments, respectively, under two control algorithms. The central line within the box indicates the median entanglement number, while the boundaries represent the 25th and 75th percentiles (lower and upper quartiles). The whiskers extend to the smallest and largest entanglement numbers within 1.5 times the interquartile range from the quartiles. The number of entanglements under the proposed TGI algorithm is notably lower than that under the BOIDS algorithm. (b) An illustration of entanglements: when parts of two or more cyborgs overlap with each other. Entanglement can lead to undesirable results, such as damage to the cyborgs. (c) A group of close but non-entangling cyborgs. (d) The minimum distance among all pairs of cyborgs circled out in experimental snapshots in one trial of experiments conducted with the BOIDS algorithm. Red points indicate instances where at least one of those cyborgs received electrical stimulation. The frequent stimulations on cyborgs that were too close to each other led to entanglements (distance less than 6 cm). (e) The minimum distance in one trial of experiments under TGI control. The TGI algorithm leverages insects’ instincts to prevent entanglements. The corresponding snapshots reveal that cyborgs can move closely without entanglements under TGI control. The experiment results are shown by Supplementary Video 3. Source data are provided as a Source Data file.
Figure 5: The process of an immobilized cyborg being saved by neighbors. (a) Entangle and detach from obstacles. A cyborg’s “Y” shape marker became wedged on an obstacle ($t=t_{2}$), and the corresponding displacement almost remained the same during this period. Then the surrounding cyborgs navigated around the trapped one and the obstacle ($t=t_{3}$). Through our algorithm, the trapped cyborg gradually overcame the obstacle with the “attractive force” from other cyborgs, and the displacement started changing again. The red dots in the displacement plot represent electrical stimulations applied to the cyborgs. Finally, the cyborg detached from the obstacle ($t=t_{4}$). See Supplementary video 4. (b) Recover from an overturn. The grey bar in the bar chart denotes the time interval of a self-attempted period for recovery of an overturned cyborg. The bar’s bottom and top indicate the start and ending moments. Similarly, the purple bar denotes the time taken for recovery with help from a neighbor. In Cases 1 and 3, the overturned cyborg first attempted and failed to recover by itself for a while, then successfully recovered by grabbing a passing neighbor. The corresponding recovery processes are illustrated by the snapshots. The experiment results are shown by Supplementary Video 5. Source data are provided as a Source Data file.